Evaluate limit x tends to 2 x to the power 10 minus 1024 divided by x x to the power 8 minus 256

Grade:11

1 Answers

Sayam
7 years ago
=lt x\rightarrow2  x10-1024/x8-256
=lt x\rightarrow2  x10-210/x8-28
=lt x\rightarrow2  (x10-210)(x-2)/(x8-28)(x-2)...........multiply and divide x-2
=lt x\rightarrow2  (x10-210)/(x-2)  X   lt x\rightarrow2  (x-2)/(x8-28)     ….....i
Applying the rule
lt x\rightarrowa  (xn-an)/(x-a)=nXan-1
now the equation i becomes-
=10X29/8X27=  5